# Solving the Poisson equation using coupled Markov chains

### v4

@@ 2.3 Interpretation and verification of Assumption 2
and $\mathbb{E}_{x,y}[\tau^{\kappa}]<\infty$ for $\pi\otimes\pi$-almost all $(x,y)$. Conversely, if for some $s>\kappa$, there exists $\tilde{C}:\mathbb{X}\times\mathbb{X}\to\mathbb{R}$ with $\pi\otimes\pi(\tilde{C})<\infty$ such that for $\pi\otimes\pi$-almost all $(x,y)$, we have
@@ 2.4 Illustration
take values within $(-a,+a)$ with $a=\max|z_{i}|$, since they are weighted averages of $(z_{i})$ multiplied by a value in $(0,1)$. Therefore, the mean of the next $\theta^{\prime}$ is within a finite interval that does not depend on the previous $\theta$. Regarding the variance $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}$, note that $\eta_{i}=-{2}/({1+(\theta-z_{i})^{2}})\log U_{i}$, where $U_{i}$ is Uniform$(0,1)$, thus $\eta_{i}\leq-2\log U_{i}$ and finally $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}\geq(\sigma^{-2}+\sum_{i=1}^{n}(-2\log U_{i}))^{-1}$. Also, $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}\leq\sigma^{2}$. The distribution of $\sum_{i=1}^{n}(-2\log U_{i})$ does not depend on $\theta$, thus we can define an interval $(c,d)\subset(0,\infty)$ and $\epsilon\in(0,1)$, independently of $\theta$ and $\tilde{\theta}$, such that $(\sum_{i=1}^{n}\eta_{i}+\sigma^{-2})^{-1}\in(c,d)$ and $(\sum_{i=1}^{n}\tilde{\eta}_{i}+\sigma^{-2})^{-1}\in(c,d)$ simultaneously with probability $\epsilon$. Therefore, with probability $\epsilon$, $(\theta^{\prime},\tilde{\theta}^{\prime})$ is drawn from a maximal coupling of two Normals, which means and variances are in finite intervals defined independently of $(\theta,\tilde{\theta})$. Two such Normals have a total variation distance that is bounded away from one, thus there exists $\delta>0$ such that, $\mathbb{P}(\theta^{\prime}=\tilde{\theta}^{\prime}|\theta,\tilde{\theta})>\epsilon\delta$, for some $\epsilon>0$, $\delta>0$ and all $(\theta$,$\tilde{\theta})$. Hence, Assumption 2 holds for all $\kappa\geq 1$.
@@ 3.3 Subsampled unbiased estimators
where $\omega_{i}$ are the weights, of the form $(\ell-k+1)^{-1}$ for the first $\ell-k+1$ atoms, and of the form ±$\pm v_{t}(\ell-k+1)^{-1}$ with $v_{t}$ defined in (3.6) and $t\in\{k+L,\ldots,\tau-1\}$ for the remaining atoms. We further simplify the notation by writing $\hat{\pi}$ instead of $\hat{\pi}_{k:\ell}$. For a function $h$, $\hat{\pi}(h)=\sum_{i=1}^{N}\omega_{i}h(Z_{i})=H_{k:\ell}$, where $H_{k:\ell}$ is the unbiased MCMC estimator in Definition 16.
@@ 4.4 Implementation and improvements
$\alpha_{n}$ =$=\{\omega_{n}(h(Z_{n})-\hat{\pi}(h))\}^{2}\mathbb{E}\left[G(Z_{n})^{2}|Z_{n}\right],$
@@
We mention other methodological variations that we do not investigate further in this manuscript. Instead of sampling $R$ atoms from each signed measure with replacement, we could sample without replacement. Also, the $2R$ estimators $G$ employed in (4.4) could be generated jointly instead of independently. In particular, we can couple $2R$ chains starting from $(Z_{I^{(j)}_{r}})_{r=1}^{R}$, $j\in\{1,2\}$ with a common chain starting from $y$: in other words, we could simulate a single coupling of $12R+$ 1 chains instead of simulating $2R$ couplings of two chains.
@@ 4.5 Multivariate extension
$-(\pi(h_{i}\cdot h_{j})-\pi(h_{i})\pi(h_{j}))+\pi\left[(h_{i}-\pi(h_{i}))g_{j}+g_{i}(h_{j}-\pi(h_{j}))\right].$ (4.9)
@@ 5.1 AR(1)
Next we compare the proposed estimator with the following naive strategy: generate a chain of length $T$ (post burn-in), compute the estimate $T^{-1}\sum_{t=0}^{T-1}h(X_{t})$, repeat $M$ times independently and compute the empirical variance of the $M$ estimates. We set $T=13,000$ to match the cost of our estimator with $R=50$. We find that the naive strategy has a bias (as $M\to\infty$) equal to $-77$, and an asymptotic variance of $2\times 10^{8}$. This is about 10 times larger than the variance of the proposed estimator reported in Table 5.1.
@@ 5.2 Hamiltonian Monte Carlo for Bayesian logistic regression
$\forall i\in[N]\quad Y_{i}\sim\text{Bernoulli with probability }1/(1+\exp(-(\alpha+x_{i\cdot}\beta+\gamma_{\psi(i)}))),$ (5.1)
@@ Acknowledgements
Research of AL supported by EPSRC grants EP/R034710/1 and EP/Y028783/1. Research of DV supported by SERB (SPG/2021/001322).
@@ B.4 Unbiased approximation of $\pi(h)$
$\mathbb{E}\left[\left|H_{k}\right|^{p}\right]^{\frac{1}{p}}\leq M^{\frac{1}{p}}\left\|h\right\|_{L^{m}(\pi)}\zeta\left(\frac{(m-p)\kappa}{mp}\right)\left\{1+2M^{\frac{m-p}{mp}}\mathbb{E}_{\pi\otimes\pi}\left[\left(0\wedge\frac{\tau-k}{L}\right)^{\kappa}\right]^{\frac{m-p}{mp}}\right\}<\infty,$ (B.9)
@@
[…]
@@
$\leq M^{\frac{1}{p}}\left\|h\right\|_{L^{m}(\pi)}\zeta\left(\frac{(m-p)\kappa}{mp}\right)\left\{1+2M^{\frac{m-p}{mp}}\mathbb{E}_{\pi\otimes\pi}\left[\left(0\wedge\frac{\tau-k}{L}\right)^{\kappa}\right]^{\frac{m-p}{mp}}\right\}$
<<$\infty$,< ∞ ,
@@
$\mathbb{E}\left[\left|H_{k:\ell}\right|^{p}\right]^{\frac{1}{p}}\leq M^{\frac{1}{p}}\left\|h\right\|_{L^{m}(\pi)}\zeta\left(\frac{(m-p)\kappa}{mp}\right)\left\{1+2M^{\frac{m-p}{mp}}\mathbb{E}_{\pi\otimes\pi}\left[\left(0\wedge\frac{\tau-k}{L}\right)^{\kappa}\right]^{\frac{m-p}{mp}}\right\}<\infty.$
@@
×$\times M^{\frac{1}{m}}\left\|h\right\|_{L^{m}(\pi)}\left\{\sum_{i=1}^{\infty}\frac{1}{i^{1+\varepsilon}}\right\}^{\frac{\kappa-p}{\kappa p}}<\infty,$
@@ B.5 Unbiased approximation of $\pi(h_{0}\cdot g)$
×$\times\left\{\mathds{1}(x^{\prime}=y^{\prime})\,\delta_{u^{\prime}}({\rm d}v^{\prime})+\mathds{1}(x^{\prime}\neq y^{\prime})\,Q({\rm d}v^{\prime})\right\}.$
@@ Appendix C Comparison with long-run variance estimators
[…]
@@ Appendix D Verifying the assumptions in the AR(1) case
A Markov chain $X$ is an ${\rm AR}(1,\phi,\sigma^{2})$ chain if for a sequence of independent $\text{Normal}(0,1)$ random variables $Z=(Z_{n})$, $X_{n}=\phi X_{n-1}+\sigma Z_{n}$ for all $n\geq 1$. We shall assume here that $\phi\in(0,1)$. We use the reflection-maximal coupling in Algorithm A.4 to construct a Markovian coupling of two AR(1,$\phi$,$\sigma^{2}$) chains $X$ and $Y$ started from $x_{0}$ and $y_{0}$, respectively. For convenience in this section, we describe the reflection-maximal coupling of $\text{Normal}(x,\sigma^{2})$ and $\text{Normal}(y,\sigma^{2})$ as follows:
@@
where $g(x,y,w)=\varphi\left(\phi\frac{(x-y)}{\sigma}+w\right)/\varphi(w)$. $X$ is the Markov chain described by the algorithm, while $Y^{\prime}$ is the Markov chain associated with the $Y$ chain when the Bernoulli random variables $B_{n}$ take the value 00 for all $n$. In particular, if $\tau=\inf\{n\geq 1:B_{n}=1\}$ then $Y_{n}=Y^{\prime}_{n}$ for $n\in\{0,\ldots,\tau-1\}$ and $Y_{n}=X_{n}$ for $n\geq\tau$. Hence, the meeting time $\tau$ can be determined by analyzing the $X$ and $Y^{\prime}$ chains, and in particular
@@
This equivalence suggests the relevance of that expectation with respect to an AR$(1,\phi,1)$ chain, which we bound below. The proof is delayed to the end of this appendix as it requires several intermediate results.
@@
where with the same constants as in Proposition D.39, $\tilde{C}(x,y)=\frac{2}{\tilde{\beta}}+|\frac{x-y}{2\sigma}|+3$ and $\bar{\beta}=\tilde{\beta}^{\frac{\log(\phi)}{\log(\tilde{\beta})+\log(\phi)}}$, and satisfies $\pi\otimes\pi(\tilde{C})<\infty$.
We see that the dependence on $\sigma$ and $|x-y|$ is fairly mild. On the other hand, if one calculates the dependence of $\bar{\beta}$ on $\phi$, one finds that it deteriorates quickly as $\phi\nearrow 1$, even though it remains less than $1$. In contrast, numerical experiments suggest that the true geometric rate is in fact $\phi$, but we are not aware of a proof technique that is able to capture such a rate. Indeed, the calculations we have used to provide a rigorous bound are similar to those used to provide quantitative convergence rates for Markov chains more generally and these are often loose in practice.
@@
$\mathbb{E}_{x}[G(x,X_{1})]$ =$=\int\left\{1-\frac{\varphi\left(2\phi x+w\right)}{\varphi(w)}\right\}_{+}\varphi(w){\rm d}w$
=$=\int\left\{\varphi(w)-\varphi\left(2\phi x+w\right)\right\}_{+}{\rm d}w$
@@
where $\tilde{\beta}\in(0,1)$ is as in Proposition D.42, and $\tilde{C}(x,y)=1+(1-\phi^{2})\left(\frac{x_{0}-y_{0}}{2\sigma}\right)^{2}$ satisfies $\pi\otimes\pi(\tilde{C})<\infty$.
@@ E.3 Coupling of HMC
    Return $(x^{\prime},\tilde{x}^{\prime})$.
@@
    Sample proposals $(x^{\star},\tilde{x}^{\star})$ from a reflection-maximal coupling of $\text{Normal}(\mu,\epsilon^{2}M^{-1})$ and $\text{Normal}(\tilde{\mu},\epsilon^{2}M^{-1})$, see Algorithm A.3.
@@
    $\tilde{x}^{\prime}$ $=\tilde{x}^{\star}\mathds{1}(U<\alpha(\tilde{x},\tilde{x}^{\star}))+\tilde{x}\mathds{1}(U\geq\alpha(\tilde{x},\tilde{x}^{\star})).$
@@
    Return $(x^{\prime},\tilde{x}^{\prime})$.
@@
    If $L=1$, obtain $(x^{\prime},\tilde{x}^{\prime})$ from the coupled MALA transition in Algorithm E.4.
@@
    If $L>1$, obtain $(x^{\prime},\tilde{x}^{\prime})$ from the coupled HMC transition in Algorithm E.3, with $L$ leapfrog steps.
@@
    Return $(x^{\prime},\tilde{x}^{\prime})$.
@@ Appendix F MCMC for Bayesian linear regression with shrinkage prior
We examine a Bayesian linear regression of $n=71$ responses on $p=4088$ predictors of the riboflavin data set (Bühlmann et al. 2014), using a continuous shrinkage prior on the coefficients (e.g. Bhadra et al. 2019). The model, MCMC algorithm and its coupling are taken from Johndrow et al. (2020), Biswas et al. (2022). Computational difficulties arise from the use of the shrinkage prior, that induces multimodality and heavy tails in the posterior distribution of the coefficients, denoted by $\beta$. The target is defined on the space of coefficients, global precision, local precisions, and the variance of the observation noise. The state space is of dimension $81782p+2=$ 8178. In this example, the meeting times have been shown to have Exponential tails in Biswas et al. (2022, Proposition 6) under the assumption that the global precision $\xi$ has a compact support, in which case Assumption 2 holds for any $\kappa$. In the experiments, we do not restrict the support of $\xi$, and we use a half-Cauchy distribution on $\xi^{-1/2}$.
@@ F.1 Model
for $\text{for }j=1,\ldots,p\quad$ $\beta_{j}\sim\text{Normal}(0,\sigma^{2}/\xi\eta_{j}),\quad\eta_{j}^{-1/2}\sim t(\nu)^{+}.$
@@ F.2.1 $\eta$-update
[…]
@@ F.2.4 $\beta$-update
Algorithm F.2 Sampling from Normal$((X^{T}X+\xi\text{Diag}(\eta))^{-1}X^{T}y,\sigma^{2}(X^{T}X+\xi\text{Diag}(\eta))^{-1})$.

### v5

@@ 2.3 Interpretation and verification of Assumption 2
[…]
@@ 2.4 Illustration
[…]
@@ 3.3 Subsampled unbiased estimators
where $\omega_{i}$ are the weights, of the form $(\ell-k+1)^{-1}$ for the first $\ell-k+1$ atoms, and of the form $\pm v_{t}(\ell-k+1)^{-1}$ with $v_{t}$ defined in (3.6) and $t\in\{k+L,\ldots,\tau-1\}$ for the remaining atoms. We further simplify the notation by writing $\hat{\pi}$ instead of $\hat{\pi}_{k:\ell}$. For a function $h$, $\hat{\pi}(h)=\sum_{i=1}^{N}\omega_{i}h(Z_{i})=H_{k:\ell}$, where $H_{k:\ell}$ is the unbiased MCMC estimator in Definition 16.
@@ 4.4 Implementation and improvements
$\alpha_{n}$ $=\{\omega_{n}(h(Z_{n})-\hat{\pi}(h))\}^{2}\mathbb{E}\left[G(Z_{n})^{2}|Z_{n}\right],$
@@
We mention other methodological variations that we do not investigate further in this manuscript. Instead of sampling $R$ atoms from each signed measure with replacement, we could sample without replacement. Also, the $2R$ estimators $G$ employed in (4.4) could be generated jointly instead of independently. In particular, we can couple $2R$ chains starting from $(Z_{I^{(j)}_{r}})_{r=1}^{R}$, $j\in\{1,2\}$ with a common chain starting from $y$: in other words, we could simulate a single coupling of $2R+1$ chains instead of simulating $2R$ couplings of two chains.
@@ 4.5 Multivariate extension
$(\pi(h_{i}\cdot h_{j})-\pi(h_{i})\pi(h_{j}))+\pi\left[(h_{i}-\pi(h_{i}))g_{j}+g_{i}(h_{j}-\pi(h_{j}))\right].$ (4.9)
@@ 5.1 AR(1)
Next we compare the proposed estimator with the following naive strategy: generate a chain of length $T$ (post burn-in), compute the estimate $T^{-1}\sum_{t=0}^{T-1}h(X_{t})$, repeat $M$ times independently and compute the empirical variance of the $M$ estimates. We set $T=13,000$ to match the cost of our estimator with $R=50$. We find that the naive strategy has a bias (as $M\to\infty$) equal to 77-77, and an asymptotic variance of $2\times 10^{8}$. This is about 10 times larger than the variance of the proposed estimator reported in Table 5.1.
@@ 5.2 Hamiltonian Monte Carlo for Bayesian logistic regression
[…]
@@ Acknowledgements
Research of AL supported by EPSRC grants EP/R034710/1 and EP/Y028783/1. Research of DV supported by SERB (SPG/2021/001322). The authors thank Mohamed Keteb for finding a typo in an earlier version of the manuscript.
@@ B.4 Unbiased approximation of $\pi(h)$
$\mathbb{E}\left[\left|H_{k}\right|^{p}\right]^{\frac{1}{p}}\leq M^{\frac{1}{p}}\left\|h\right\|_{L^{m}(\pi)}\zeta\left(\frac{(m-p)\kappa}{mp}\right)\left\{1+2M^{\frac{m-p}{mp}}\mathbb{E}_{\pi\otimes\pi}\left[\left(0\vee\left\lceil\frac{\tau-k}{L}\right\rceil\right)^{\kappa}\right]^{\frac{m-p}{mp}}\right\}<\infty,$ (B.9)
@@
$\tau_{L,k}\overset{d}{=}0\vee\left\lceil\frac{\tau-k}{L}\right\rceil,$
@@
$\leq M^{\frac{1}{p}}\left\|h\right\|_{L^{m}(\pi)}\zeta\left(\frac{(m-p)\kappa}{mp}\right)\left\{1+2M^{\frac{m-p}{mp}}\mathbb{E}_{\pi\otimes\pi}\left[\left(0\vee\left\lceil\frac{\tau-k}{L}\right\rceil\right)^{\kappa}\right]^{\frac{m-p}{mp}}\right\}$
$<\infty,$
@@
$\mathbb{E}\left[\left|H_{k:\ell}\right|^{p}\right]^{\frac{1}{p}}\leq M^{\frac{1}{p}}\left\|h\right\|_{L^{m}(\pi)}\zeta\left(\frac{(m-p)\kappa}{mp}\right)\left\{1+2M^{\frac{m-p}{mp}}\mathbb{E}_{\pi\otimes\pi}\left[\left(0\vee\left\lceil\frac{\tau-k}{L}\right\rceil\right)^{\kappa}\right]^{\frac{m-p}{mp}}\right\}<\infty.$
@@
$\times M^{\frac{1}{m}}\left\|h\right\|_{L^{m}(\pi)}\left\{\sum_{i=1}^{\infty}\frac{1}{i^{1+\varepsilon}}\right\}^{\frac{\kappa-p}{\kappa p}}<\infty,$
@@ B.5 Unbiased approximation of $\pi(h_{0}\cdot g)$
$\times\left\{\mathds{1}(x^{\prime}=y^{\prime})\,\delta_{u^{\prime}}({\rm d}v^{\prime})+\mathds{1}(x^{\prime}\neq y^{\prime})\,Q({\rm d}v^{\prime})\right\}.$
@@ Appendix C Comparison with long-run variance estimators
$\sqrt{a_{t}}\left(v^{\text{BM}}(P,h)-v(P,h)\right)\overset{d}{\to}\text{Normal}(0,2v(P,h)^{2}).$
@@ Appendix D Verifying the assumptions in the AR(1) case
[…]
@@
where $g(x,y,w)=\varphi\left(\phi\frac{(x-y)}{\sigma}+w\right)/\varphi(w)$. $X$ is the Markov chain described by the algorithm, while $Y^{\prime}$ is the Markov chain associated with the $Y$ chain when the Bernoulli random variables $B_{n}$ take the value 0 for all $n$. In particular, if $\tau=\inf\{n\geq 1:B_{n}=1\}$ then $Y_{n}=Y^{\prime}_{n}$ for $n\in\{0,\ldots,\tau-1\}$ and $Y_{n}=X_{n}$ for $n\geq\tau$. Hence, the meeting time $\tau$ can be determined by analyzing the $X$ and $Y^{\prime}$ chains, and in particular
@@
[…]
@@
[…]
We see that the dependence on $\sigma$ and |xy||x-y| is fairly mild. On the other hand, if one calculates the dependence of $\bar{\beta}$ on $\phi$, one finds that it deteriorates quickly as $\phi\nearrow 1$, even though it remains less than $1$. In contrast, numerical experiments suggest that the true geometric rate is in fact $\phi$, but we are not aware of a proof technique that is able to capture such a rate. Indeed, the calculations we have used to provide a rigorous bound are similar to those used to provide quantitative convergence rates for Markov chains more generally and these are often loose in practice.
@@
$\mathbb{E}_{x}[G(x,X_{1})]$ $=\int\left\{1-\frac{\varphi\left(2\phi x+w\right)}{\varphi(w)}\right\}_{+}\varphi(w){\rm d}w$
$=\int\left\{\varphi(w)-\varphi\left(2\phi x+w\right)\right\}_{+}{\rm d}w$
@@
[…]
@@ E.3 Coupling of HMC
    Return (x,x~)($^{\prime},\tilde{x}^{\prime}$).
@@
    Sample proposals (x,x~)($^{\star},\tilde{x}^{\star}$) from a reflection-maximal coupling of $\text{Normal}(\mu,\epsilon^{2}M^{-1})$ and $\text{Normal}(\tilde{\mu},\epsilon^{2}M^{-1})$, see Algorithm A.3.
@@
    $\tilde{x}^{\prime}$ $).\displaystyle=\tilde{x}^{\star}\mathds{1}(U<\alpha(\tilde{x},\tilde{x}^{\star}))+\tilde{x}\mathds{1}(U\geq\alpha(\tilde{x},\tilde{x}^{\star}$)).
@@
    Return (x,x~)($^{\prime},\tilde{x}^{\prime}$).
@@
    If $L=1$, obtain (x,x~)($^{\prime},\tilde{x}^{\prime}$) from the coupled MALA transition in Algorithm E.4.
@@
    If $L>1$, obtain (x,x~)($^{\prime},\tilde{x}^{\prime}$) from the coupled HMC transition in Algorithm E.3, with $L$ leapfrog steps.
@@
    Return (x,x~)($^{\prime},\tilde{x}^{\prime}$).
@@ Appendix F MCMC for Bayesian linear regression with shrinkage prior
We examine a Bayesian linear regression of $n=71$ responses on $p=4088$ predictors of the riboflavin data set (Bühlmann et al. 2014), using a continuous shrinkage prior on the coefficients (e.g. Bhadra et al. 2019). The model, MCMC algorithm and its coupling are taken from Johndrow et al. (2020), Biswas et al. (2022). Computational difficulties arise from the use of the shrinkage prior, that induces multimodality and heavy tails in the posterior distribution of the coefficients, denoted by $\beta$. The target is defined on the space of coefficients, global precision, local precisions, and the variance of the observation noise. The state space is of dimension $2p+2=8178$. In this example, the meeting times have been shown to have Exponential tails in Biswas et al. (2022, Proposition 6) under the assumption that the global precision $\xi$ has a compact support, in which case Assumption 2 holds for any $\kappa$. In the experiments, we do not restrict the support of $\xi$, and we use a half-Cauchy distribution on $\xi^{-1/2}$.
@@ F.1 Model
$\text{for }j=1,\ldots,p\quad$ $\beta_{j}\sim\text{Normal}(0,\sigma^{2}/\xi\eta_{j}),\quad\eta_{j}^{-1/2}\sim t(\nu)^{+}.$
@@ F.2.1 $\eta$-update
    where $\gamma_{s}(x):=\Gamma(s)^{-1}\int_{0}^{x}t^{s-1}e^{-t}dt\in[0,1]$ is the cdf of the $\mathrm{Gamma}(s,1)$ distribution.
@@ F.2.4 $\beta$-update
[…]